During a training session a footballer kicked a ball with a velocity of 14ms-1 and another footballer kicked a second ball of equal mass in the opposite direction with a velocity of 20ms-1. The two balls made a head on head collision and the first ball began to move with a velocity of 16ms-1 in the opposite direction find the new velocity of the second ball each ball is 900g
1
Expert's answer
2020-02-25T10:26:03-0500

From the conservation of momentum:


"mv_1-mu_1=-mv_2+mu_2"

"v_1-u_1=-v_2+u_2"

"14-20=-16+u_2"

"u_2=10\frac{m}{s}"